f(x)=cospx+sinpx is periodic then p must be rational or irrational or can be postive real number.
Thanks

Grade:

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
f(x) = cos(x)
f(x) = sin(x)
are periodic function with a period of 2pi.
f(x) = cos(px)
f(x) = sin(px)
are periodic with a period of
\frac{2\pi }{|p|}, p\neq 0
p can be anything other than zero.
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty